OpenStudy (anonymous):

Find the inverse of f(x)=x^3+2

jimthompson5910 (jim_thompson5910):

\[\large f(x)=x^3+2\] \[\large y=x^3+2\] \[\large x=y^3+2\] \[\large x-2=y^3\] \[\large \sqrt[3]{x-2}=y\] \[\large y=\sqrt[3]{x-2}\] Inverse is \[\large f^{-1}(x)=\sqrt[3]{x-2}\]